OpenStudy (lilthesis101):

A triangle has side lengths 10, 15, and 7. Is the triangle a cute, obtuse, or right ? Show your work and explain.

OpenStudy (lilah):

hey try tagging people

OpenStudy (lilthesis101):

What do you mean?

jimthompson5910 (jim_thompson5910):

we have a triangle with sides a,b,c a = 10 b = 7 c = 15 make 'c' the longest side. The 'a' and 'b' can be in either order. It doesn't matter now tell me what a^2+b^2 = 10^2+7^2 is equal to

OpenStudy (lilthesis101):

149

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

what is c^2 = 15^2 equal to?

OpenStudy (lilthesis101):

225

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

now compare the two results a^2+b^2 = 149 c^2 = 225 we see that the `c^2` expression is larger, so \(\Large c^2 > a^2+b^2\) . Based on these rules here http://images.slideplayer.com/8/2473263/slides/slide_4.jpg this means that we have an obtuse triangle